Determine the limit of (1+2+...+n)/(n^2) when n goes to infinity.
Can I solve this by taking the sequence to n brackets?
1
Expert's answer
2020-12-21T18:35:45-0500

Let "a_n=\dfrac{1+2+...+n}{n^2}." Then


"a_n=\dfrac{1+2+...+n}{n^2}"

"a_n=\dfrac{\displaystyle\bigg(\sum_{i=1}^i\bigg)}{n^2}"

"a_n=\dfrac{\displaystyle\dfrac{n(n+1)}{2}}{n^2}"

"a_n=\dfrac{n^2+n}{2n^2}"

"a_n=\dfrac{1}{2}+\dfrac{1}{2n}"

"\lim\limits_{n\to \infin}\dfrac{1+2+...+n}{n^2}=\lim\limits_{n\to \infin}a_n"

"=\lim\limits_{n\to \infin}\big(\dfrac{1}{2}+\dfrac{1}{2n}\big)=\dfrac{1}{2}+0"

"=\dfrac{1}{2}"

Therefore


"\lim\limits_{n\to \infin}\dfrac{1+2+...+n}{n^2}=\dfrac{1}{2}"